Consider the following table

Called in Sick Went to Work Total
Had the flu 11 1 12
Was not sick 2 102

104

Total 13 103 116

If we randomly select an employee, determine the probability of selecting

a) Someone with the flu

b) A person that came to work, even though they had the flu.

c) A person that was not sick, but called in sick.

d) A person that wasn’t sick and came to work.

e) CfcvwihDK+KQP9fCw3cy511YZvHOxpCXx2OeiME3A person that wasn’t sick or came to work.


.
Answer #1

a)

probability of selecting someone with the flu = Number of people with the flu / Total Number of people

= 12 / 116 = 0.1034483

b)

probability of a person that came to work, even though they had the flu = Number of people that came to work, even though they had the flu/ Total Number of people

= 1/ 116 = 0.00862069

c)

probability of person that was not sick, but called in sick = Number of people was not sick,and called in sick / Total Number of people = 2/116 = 0.01724138

d)

probability of person that wasn’t sick and came to work = Number of people that wasn’t sick and came to work/ Total Number of people = 102 / 116 = 0.8793103